OpenStudy (anonymous):

Solve the system:3x-2y= 10 -6x+5y= -2 write your answer as an ordered pair (x,y). leave answers in fraction form if needed

OpenStudy (anonymous):

(equation one)\[=3x-2y =10\] \[2(3x-2y)=2(10)\] \[6x-4y=2\] *all i did was multiply both sides by 2 to get 6x in the equation to cancel when adding equation one and two together (equation two) \[-6x+5y=-2)\] ADD BOTH EQUATIONS(add left sides up and right sides up) \[6x-4y+(-6x+5y)=20+(-2)\] \[6x-4y-6x+5y=20-2\] \[6x-6x+5y-4y=18\] \[(6-6)x+(5-4)y=18\] \[0x+y=18\] \[y=18\] take your y answer and put it into one of the equations you added together \[-6x+5y=-2\] \[-6x+5(18)=-2\] \[-6x=-92\] \[x=\frac{-92}{-6}=\frac{92}{6}\]

OpenStudy (anonymous):

can you simplify 6/92?

OpenStudy (anonymous):

\[\frac{92}{6}=\frac{2}{2}*\frac{46}{3}\]

OpenStudy (anonymous):

\[=1*\frac{46}{3}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

btw @Trplingup i think you forgot to divide the right side when you divided by two... but you deleted before i could check =/

OpenStudy (anonymous):

Yea I did I deleted since it was wrong too bad you can't edit